User Tools

Site Tools


This wiki is not maintained! Do not use this when setting up AuScope experiments!



Module: HOB+0108, +0GB

  • Good pointing fit. (Dave)
  • 22:10 UT: Autocorrelation spectra steep in bands 07 and 08 (inverted) and 10, and ludicrously steep in bands 15 and 16 (Ross).


Module: HOB+1010, +0GB

  • Could not get a good pointing fit. (Dave)


Module: HOB+1007, +0GB

  • Good pointing fit. (Dave)
  • eRc refusing to connect. Using backup FS interface. (Dave)
  • Tsys still looking good! AGC in IFC is close to saturated though at ~62. (Dave)
  • 22:05 UT: eRemote Control restarted and working fine (Ross).
You could leave a comment if you were logged in.
/home/www/auscope/opswiki/data/pages/handover/aov001.txt · Last modified: 2015/03/21 22:10 by David Horsley